# Multi-Level Querying using A Knowledge Pyramid

### v2

@@ Abstract
This paper addresses the need for improved precision in existing Retrieval-Augmented Generation (RAG) methods that primarily focus on enhancing recall. We propose a multi-layer knowledge pyramid approach within the RAG framework to achieve a better balance between precision and recall. The knowledge pyramid consists of three layers: Ontologies, Knowledge Graphs (KGs), and chunk-based raw text. We employ cross-layer augmentation techniques for comprehensive knowledge coverage and dynamic updates of the Ontology schema and instances. To ensure compactness, we utilize cross-layer filtering methods for knowledge condensation in KGs. Our approach, named PolyRAG, follows a waterfall model for retrieval, starting from the top of the pyramid and progressing down until a confident answer is obtained. We introduce two benchmarks for domain-specific knowledge retrieval, one in the academic domain and the other in the financial domain. The effectiveness of the methods has been validated through comprehensive experiments by outperforming 19 SOTA methods. An encouraging observation is that the proposed method has augmented the GPT-4, providing 395% F1 gain by improving its performance from 0.1636 to 0.8109.
Multi-Level Querying using A Knowledge Pyramid
Rubing Chen1, Xulu Zhang1, Jiaxin Wu1, Wenqi Fan1, Xiao-Yong Wei 1, Qing Li1,* 1The Hong Kong Polytechnic University
1 Introduction
@@
Figure 1: PolyRAG Framework based on the concept of Knowledge Pyramid: different layers signify varying levels of specificity and information richness. The layers are mutually complementary, working together to enhance the overall retrieval and generation process.
@@
Significant advancements have been made in Large Language Models (LLMs), including proprietary models like ChatGPT and GPT-4, as well as open-source variants like FLAN Wei et al. (2021) and LLaMA Touvron et al. (2023a). These models have demonstrated remarkable achievements across a wide range of general knowledge tasks in areas such as language comprehension Radford et al. (2018); Hadi et al. (2023), logical reasoning Kojima et al. (2022); Chang et al. (2024), and complex question-answering Wu et al. (2023b). However, the one-size-fits-all nature of general LLMs fails to meet the specific demands for professional or personalized knowledge, such as law Huang et al. (2023b); Cui et al. (2023) and finance domains Wang et al. (2023); Xie et al. (2023). A straightforward solution is to utilize Supervised Fine-Tuning (SFT) to tailor LLMs to domain-specific tasks Ouyang et al. (2022). Examples include AlpaCare Zhang et al. (2023), Mental LLaMA Yang et al. (2024a) and Zhongjing Yang et al. (2024b) in medical domain, BloombergGPT Wu et al. (2023a), Pixiu Xie et al. (2023), and DISC-FinLLM Chen et al. (2023) in financial domain, and DISC-LawLLM Yue et al. (2023) in law domain. Nonetheless, this risks catastrophic forgetting Luo et al. (2023) of the general knowledge when tuning the original model, and is prone to model hallucination Huang et al. (2023a).
@@
To address this challenge, the prevalent alternative of Retrieval Augmented Generation (RAG) is introduced as a means to enhance the domain knowledge comprehension of LLMs Lewis et al. (2020a). Rather than solely relying on the generation capabilities of LLMs to produce answers, RAG takes a different approach by incorporating information retrieval techniques. It retrieves relevant information from existing resources and utilizes this information to enrich the context of the prompt Gao et al. (2023). This enables in-context learning Dong et al. (2023) or few-shot learning Wang et al. (2020b) to be initiated based on the enhanced context. It makes the LLMs’ output more stable, accurate, traceable, and interpretable. However, early implementations of RAG have relied on unstructured textual data chunks, which are often obtained by partitioning each document in the corpora database into segments with a predefined chunk size and thus not organized in a specific way Ma et al. (2023); Lewis et al. (2020b). To address this limitation, Knowledge Graphs (KGs) have been incorporated into RAG Baek et al. (2023a); Wu et al. (2023d); Abu-Rasheed et al. (2024). The incorporation of KGs in RAG is a logical choice, as KGs have long been recognized as a promising augmentation method for retrieval tasks. In fact, in traditional retrieval tasks, another solution called Ontologies has been found to be suitable for domain-specific tasks due to their formally and strictly defined schemas Vamsi et al. (2024); Palagin et al. (2023). However, the use of Ontologies in RAG has been underexplored, partly because the process of defining the Ontology schema requires significant human effort.
@@
The observation serves as motivation for us to build a multi-layer knowledge pyramid and devise a customized querying strategy within the RAG framework. The main objective of this approach is to achieve a better balance between “precision” and “recall”. Specifically, we organize a knowledge pyramid with three distinct layers of Ontologies, knowledge graphs, and chunk-based raw text. During the construction of the knowledge pyramid, we ensure comprehensive coverage of knowledge by incorporating cross-layer augmentation techniques. This approach reduces the need for human intervention in knowledge completion, as it allows for dynamic updates of the Ontology schema and instances. Moreover, we maintain compactness by employing cross-layer filtering methods, which remove the redundant triplets in the KG for knowledge condensation. The retrieval process follows a waterfall model, starting from the top of the pyramid and progressively moving down until the model reaches a confident answer. If certainty is not achieved, the search continues to the next layer down, and so on. We name it as PolyRAG due to its nature of multi-layer knowledge organization and multi-level querying strategy.
@@
Furthermore, this paper presents a contribution by introducing two benchmarks for domain-specific knowledge retrieval. The first benchmark is specifically designed for the academic domain and has been meticulously constructed by the authors themselves. It encompasses extensive information regarding 1,319 staff members, 2,061 courses, 31 departments, as well as classrooms and facilities at XXX university. In addition, the authors have extended an existing public dataset called FiQA Shah et al. (2022) in the financial domain to create the second benchmark. We restructured the knowledge from this dataset into a similar pyramid structure, allowing for the application of the proposed methodology. Both benchmarks will be made available to the community.
2 Related Works
2.1 Domain-Specific Large Language Models
@@
LLMs have become integral in various applications, such as chatbots, writing assistants, customer service automation Sallam (2023); Rebedea et al. (2023). Domain-specific LLMs are a subset of LLMs that have been tailored to understand and generate text within a particular area of expertise, such as law Huang et al. (2023b); Cui et al. (2023); Yue et al. (2023), medicine Zhang et al. (2023); Yang et al. (2024b, a), or finance Ge et al. (2024); Wang et al. (2023); Xie et al. (2023); Chen et al. (2023). These models aim to provide higher accuracy and more relevant outputs than general-purpose LLMs when dealing with specialized content Ouyang et al. (2022). A common approach is to adapt instruction fine-tuning based on the general LLMs Ouyang et al. (2022) using domain-specific copra. However, research has shown that the generality of the fine-tuned model will diminish, resulting in catastrophic forgetting Luo et al. (2023); Zhai et al. (2023). In order to tackle the limitations, the approach of prompting LLMs with extra domain knowledge as contexts is widely adopted.
2.2 Retrieval-Augmented Generation
@@
Retrieval-Augmented Generation (RAG) enhances the generative capabilities of language models by incorporating retrieved knowledge for in-context learning Lewis et al. (2020b); Gao et al. (2023). NaiveRAG represents the most basic architecture within this framework, in which the system retrieves the top-k documents that are most relevant to the query and integrate them into the prompt, thereby grounding the responses in more relevant information Ma et al. (2023).
@@
Expanding on NaiveRAG, advanced RAG incorporates additional modules or structures to improve retrieval precision. Reranking is a notable example, where a reranker is employed to refine the initial ranked list (e.g., Re2G Glass et al. (2022) and bge-reranker Xiao et al. (2023), both are based on BERT Kenton and Toutanova (2019)). Furthermore, studies have indicated that excessive noise and lengthy context can have a negative impact on inference performance. To address this, prompt compression methods such as Selective Context Litman et al. (2020) and LLMLingua Anderson et al. (2022) have been developed. These methods emphasize key information while reducing noise and context length, as discussed in Gao et al. (2023).
2.3 Knowledge-Augmented Language Models
@@
The Knowledge-Augmented Language Model (LM) approach involves integrating LMs with additional knowledge bases to facilitate in-context learning Liu et al. (2019). In addition to incorporating knowledge in the form of raw texts Ma et al. (2023), knowledge graphs (KGs) Baek et al. (2023b); Pan et al. (2024); Wu et al. (2023d) have gained popularity, and Ontologies Vamsi et al. (2024) have also been utilized, albeit less frequently. In most KG augmentation methods Pan et al. (2024), the RAG framework is followed, wherein the context is expanded by incorporating retrieved KG triples instead of raw text chunks. KAPING Baek et al. (2023a) serves as an early example of this approach, which has been later refined in RRA Wu et al. (2023c). Regarding Ontologies, instead of being used as an individual knowledge base, they are often employed as assistants for generating KG triplets Pan et al. (2024) (e.g., Text2KGBench Mihindukulasooriya et al. (2023)) or for augmenting textual corpora (e.g., EKGs Baldazzi et al. (2023), OntoChatGPT Palagin et al. (2023)). The integration of various forms of knowledge bases poses a significant challenge and remains an area that has received limited exploration.
3 Knowledge Pyramid Construction
@@
3.1 Construction of Initial Layers
Ontology Layer: The Ontology layer $\mathcal{O}=\{\mathcal{O}_{s},\mathcal{O}_{i}\}$ consists of a schema $\mathcal{O}_{s}$ and corresponding instances $\mathcal{O}_{i}$. Defining an Ontology schema typically requires significant time and effort from human experts, and achieving a comprehensive schema can be challenging. To simplify the process, one approach is to initially extract a sub-domain schema from general Ontologies like WordNet Miller (1995) or ConceptNet Speer et al. (2017). This extracted schema can serve as a starting point and be refined using the semi-automatic approach presented in Section 3.2.
With the schema $\mathcal{O}_{s}$ that we extracted, we can guide LLMs to extract instances from the raw text layer $\mathcal{T}$ for each of the concept-attribute pair $(c,a)\in\mathcal{O}_{s}$, where $c$ is a concept (e.g., professor) and $a$ is one of its attributes (e.g., research_interest). This can be written as a prompt function
$f_{ins}(c,a;\,p)$: Given a paragraph {p} from the {domain} domain, please identify instances of the Ontology relationship where a class {c} has the attribute of {a}. Note that the attribute may consist of multiple entities.
@@ 3.1 Construction of Initial Layers
Knowledge Graph Layer: We adopt Open Information Extraction (OpenIE) Etzioni et al. (2008) to extract KG triples from raw texts. However, direct extraction often results in noisy output, including irrelevant or duplicate entities. To address this, we draw inspiration from the multi-round prompting approach of LLM2KG Carta et al. (2023) and reimplement it by introducing four functions: $f_{par}(p)$ for paraphrasing, $f_{ent}()$ for entity extraction, $f_{rel}()$ for relation completion, and $f_{dis}()$ for disambiguation. These functions form a cascade for KG triplet extraction and refinement as
@@
Due to space limitations, the implementations of these functions are provided in the Appendix A.
3.2 Knowledge Completion
As shown in the pyramid of Figure 1, the higher layers are more structured but not easy to define or extract, resulting in limited coverage. One common issue is the absence of important classes or attributes from the expert-defined Ontology schema, which has historically hindered the ease of defining Ontologies. By utilizing the pyramid framework, we can address this issue in a data-driven manner. The idea is to identify noteworthy concepts and relations that exist in the lower layers but are absent from the higher layers. These identified elements are then incorporated into the higher layers to enhance knowledge completion. For instance, our research reveals that the relationship publications_in_important_journals appears frequently in both the knowledge graph and raw texts but is absent from the Ontology. This is due to the tendency of human-defined schema to overlook this specific relationship, as the presence of the publications attribute may make it seem redundant. However, users often query this specific relation, and it is explicitly mentioned in professors’ profiles. It is thus worth including it in the Ontology.
Our approach involves modeling the semantic distributions of the Ontology and knowledge graph to identify concepts and relations that exhibit significant divergence between the two layers. To achieve this, we begin by transforming class-attribute pairs in the Ontology layer $\mathcal{O}$ into subject+relation phrases that align with the format in the knowledge graph layer $\mathcal{K}$. Instructor embedding Su et al. (2022) of phrases at both layers are then encoded so as to project them into a common semantic space. We then learn for each layer a multivariate Gaussian using

### v3

# Knowledge Pyramid Construction for Multi-Level Retrieval-Augmented Generation

@@ Abstract
This paper addresses the need for improved precision in existing knowledge-enhanced question-answering frameworks, specifically Retrieval-Augmented Generation (RAG) methods that primarily focus on enhancing recall. We propose a multi-layer knowledge pyramid approach within the RAG framework to achieve a better balance between precision and recall. The knowledge pyramid consists of three layers: Ontologies, Knowledge Graphs (KGs), and chunk-based raw text. We employ cross-layer augmentation techniques for comprehensive knowledge coverage and dynamic updates of the Ontology schema and instances. To ensure compactness, we utilize cross-layer filtering methods for knowledge condensation in KGs. Our approach, named PolyRAG, follows a waterfall model for retrieval, starting from the top of the pyramid and progressing down until a confident answer is obtained. We introduce two benchmarks for domain-specific knowledge retrieval, one in the academic domain and the other in the financial domain. The effectiveness of the methods has been validated through comprehensive experiments by outperforming 19 SOTA methods. An encouraging observation is that the proposed method has augmented the GPT-4, providing 395% F1 gain by improving its performance from 0.1636 to 0.8109.
Index Terms:
Knowledge Management, Language models, Knowledge retrieval
I Introduction
@@
Figure 1: PolyRAG Framework based on the concept of Knowledge Pyramid: different layers signify varying levels of specificity and richness. The layers are mutually complementary, working together for a balance between recall and precision.
@@
Significant advancements have been made in Large Language Models (LLMs), including proprietary ones like ChatGPT and GPT-4, as well as open-source variants like FLAN [1] and LLaMA [2]. These models have demonstrated remarkable achievements across a wide range of general knowledge tasks such as language comprehension [3, 4], logical reasoning [5, 6], and complex question-answering [7]. However, the one-size-fits-all nature of general LLMs fails to meet the specific demands for professional or personalized knowledge, such as law [8, 9] and finance domains [10, 11]. A straightforward solution is to utilize Supervised Fine-Tuning (SFT) to tailor LLMs to domain-specific tasks [12]. Examples include AlpaCare [13], Mental LLaMA [14] and Zhongjing [15] in medical domain, BloombergGPT [16], Pixiu [11], and DISC-FinLLM [17] in financial domain, and DISC-LawLLM [18] in law domain. Nonetheless, this risks catastrophic forgetting [19] of the general knowledge, and is prone to model hallucination [20].
@@
Prevalent alternatives of Retrieval Augmented Generation (RAG) are thus introduced as a means to enhance the domain-specific knowledge comprehension [21]. Rather than solely relying on the generation capabilities of LLMs to produce answers, RAG takes a different approach by incorporating information retrieval techniques. It retrieves relevant information from existing resources and utilizes this information to enrich the context of the prompt [22]. This enables in-context learning [23] or few-shot learning [24] and makes the LLMs’ output more stable, accurate, traceable, and interpretable. However, early implementations of RAG have relied on unstructured textual data chunks, which are often obtained by partitioning each document in the corpora database into segments with a predefined chunk size and thus not organized in a specific way [25, 26]. To address this limitation, Knowledge Graphs (KGs) have been incorporated into RAG [27, 28, 29]. The incorporation of KGs in RAG is a logical choice, as KGs have long been recognized as a promising augmentation method for retrieval tasks. In fact, in traditional retrieval tasks, another solution called Ontologies has been found to be suitable for domain-specific tasks due to their formally and strictly defined schemas [30, 31]. Nonetheless, the use of Ontologies in RAG has been underexplored, partly because the process of defining the Ontology schema requires significant human effort.
@@
The observations we have made throughout our research motivate us to develop a multi-layer knowledge pyramid and design a customized querying strategy within the RAG framework. The main objective of this approach is to achieve a better balance between “precision” and “recall” when retrieving information from heterogeneous knowledge sources. By organizing knowledge into hierarchical layers, we can leverage the strengths of each type of knowledge representation to enhance overall system performance in information retrieval and question-answering tasks. Specifically, we structure the knowledge pyramid with three distinct layers: Ontologies, Knowledge Graphs (KGs), and chunk-based raw text. The Ontology layer provides a highly structured semantic framework with formal definitions and relationships curated by domain experts, enhancing precision. The KG layer captures entities and their interrelations extracted from data, offering a rich semantic network that balances precision and recall. The raw text layer comprises unstructured textual data, providing extensive coverage and improving recall.
Moreover, during the construction of this pyramid, we employ cross-layer augmentation techniques to ensure comprehensive coverage, dynamically updating the Ontology schema and instances without extensive human intervention. We also maintain compactness by applying cross-layer filtering methods, removing redundant triplets in the KG layer for knowledge condensation. This method enables the layers to interact naturally through the guidance of the distribution in semantic space, which has proven to be effective for creating a more comprehensive and compact knowledge base.
The retrieval process follows a waterfall model, starting from the top of the pyramid and progressively moving down until a confident answer is reached. The system first attempts to resolve queries using the most structured and precise knowledge in the Ontology layer. If certainty is not achieved, the search continues to the KG layer, and then to the raw text layer if necessary. This strategy ensures efficient retrieval by prioritizing higher-precision sources while still accessing broader information when needed. We name our approach PolyRAG to reflect its nature of multi-layer knowledge organization and multi-level querying strategy, signifying the incorporation of knowledge from multiple layers and the execution of multi-level querying during the search process. The overall framework of PolyRAG is shown in Figure 2.
@@
Furthermore, this paper presents a contribution by introducing two benchmarks for domain-specific knowledge retrieval. The first benchmark is specifically designed for the academic domain and has been meticulously constructed by the authors themselves. It encompasses extensive information regarding 1,319 staff members, 2,061 courses, 31 departments, as well as classrooms and facilities at XXX University. In addition, the authors have extended an existing public dataset called FiQA [32] in the financial domain to create the second benchmark. We restructured the knowledge from this dataset into a similar pyramid structure, allowing for the application of the proposed methodology. Both benchmarks will be made available to the community.
Figure 2: The framework of PolyRAG includes: (1) Knowledge Pyramid Construction, starting from extracting knowledge from corpora data to form the initial layers, then performing the layer interactions, which are knowledge completion and condensation. (2) Given the refined Knowledge Pyramid, the LLM inference obtains the contexts through a multi-level query, which retrieves the knowledge from the higher layer down to the lower layer in a waterfall pattern.
II Related Works
II-A Domain-Specific Large Language Models
@@
LLMs have become integral in various applications, such as chatbots, writing assistants, customer service automation [33, 34]. Domain-specific LLMs are a subset of LLMs that have been tailored to understand and generate text within a particular area of expertise, such as law [8, 9, 18], medicine [13, 15, 14], or finance [35, 10, 11, 17]. These models aim to provide higher accuracy and more relevant outputs than general-purpose LLMs when dealing with specialized content [12]. A common approach is to adapt instruction fine-tuning based on the general LLMs [12] using domain-specific copra. However, research has shown that the generality of the fine-tuned model will diminish, resulting in catastrophic forgetting [19, 36]. To tackle the limitations, the approach of prompting LLMs with extra domain knowledge as contexts is widely adopted.
II-B Retrieval-Augmented Generation
@@
Retrieval-Augmented Generation (RAG) enhances the generative capabilities of language models by incorporating retrieved knowledge for in-context learning [26, 22]. While general language models excel in producing responses for general queries, they are prone to generating hallucinations when tasked with domain-specific knowledge usage. RAG addresses this issue by retrieving established knowledge corpora and providing this information as context to the language model [22]. NaiveRAG represents the most basic architecture within this framework, in which the system retrieves the top-k documents that are most relevant to the query and integrate them into the prompt, thereby grounding the responses in more relevant information [25].
@@
Expanding on NaiveRAG, advanced RAG incorporates additional modules or structures to improve retrieval precision. Reranking is a notable example, where a reranker is employed to refine the initial ranked list (e.g., Re2G [37] and bge-reranker [38], both are based on BERT [39]). Furthermore, studies have indicated that excessive noise and lengthy context can have a negative impact on inference performance. To address this, prompt compression methods such as Selective Context [40] and LLMLingua [41] have been developed. These methods emphasize key information while reducing noise and context length, as discussed in [22]. While current knowledge retrieval methods that rely on single queries can offer some contextual assistance for domain-specific knowledge, they heavily rely on the expressive abilities of the original collections of information and consistently require a trade-off between more detailed knowledge and less noisy information [42, 43]. These techniques do not effectively tackle the issues of retrieving responses that require integrating information from numerous sources, nor do they fulfill the need for dense knowledge in question-and-answer interactions. To address these constraints, we suggest employing varied knowledge representations. Our strategy seeks to optimize the distribution of query duties among knowledge bases in a knowledge pyramid, with the goal of reducing redundancy and increasing the amount and accuracy of information.
II-C Knowledge-Augmented Language Models
@@
The Knowledge-Augmented Language Model approach involves integrating LMs with additional knowledge bases to facilitate in-context learning [44]. In addition to incorporating knowledge as raw texts [25], knowledge graphs (KGs) [45, 46, 28] have gained popularity, and Ontologies [30] have also been utilized, albeit less frequently. In most KG augmentation methods [46], the RAG framework is followed by replacing the raw text chunks with retrieved KG triplets. KAPING [27] serves as an early example of this approach, which has been later refined in RRA [47]. Regarding Ontologies, instead of being used as an individual knowledge base, they are often employed as assistants for generating KG triplets [46] (e.g., Text2KGBench [48]) or for augmenting textual corpora (e.g., EKGs [49], OntoChatGPT [31]). The integration of various forms of knowledge bases poses a significant challenge.
III Knowledge Pyramid Construction
@@
III-A Construction of Initial Layers
Ontology Layer: The Ontology layer $\mathcal{O}=\{\mathcal{O}_{s},\mathcal{O}_{i}\}$ consists of a schema $\mathcal{O}_{s}$ and corresponding instances $\mathcal{O}_{i}$. Defining an Ontology schema typically requires significant time and effort from human experts, and achieving a comprehensive schema can be challenging. To simplify the process, one approach is to initially extract a sub-domain schema from general Ontologies like WordNet [50] or ConceptNet [51]. This extracted schema can serve as a starting point and be refined using the semi-automatic approach presented in the Knowledge Completion.
With the schema $\mathcal{O}_{s}$ that we extracted using Web Ontology Language (OWL) [52], we can guide LLMs to extract instances from the raw text layer $\mathcal{T}$ for each of the class-property pair $(c,a)\in\mathcal{O}_{s}$, where $c$ is a class (e.g., staff) and $a$ is one of its property (e.g., graduated_from). This can be written as a prompt function
$f_{ins}(c,a;\,p)$: Given a paragraph {p} from the {domain} domain, please identify instances of the Ontology relationship where a class {c} has the property of {a}. Note that the property may consist of multiple entities.
@@ III-A Construction of Initial Layers
Figure 3: Knowledge completion through the cross-layer interactions. The comprehensiveness of the initial Ontology schema has been improved with the noteworthy concepts identified from the knowledge graph layer.
Knowledge Graph Layer: We adopt Open Information Extraction (OpenIE) [53] to extract KG triplets from raw texts. However, direct extraction often results in noisy output, including irrelevant or duplicate entities. To address this, we draw inspiration from the multi-round prompting approach of LLM2KG [54] and reimplement it by introducing four functions: $f_{par}(p)$ for paraphrasing, $f_{ent}()$ for entity extraction, $f_{rel}()$ for relation completion, and $f_{dis}()$ for disambiguation. These functions form a cascade for KG triplet extraction and refinement as
@@
Specifically, the detailed implementation of four functions is shown as
Therefore, the initial layers, including both ontology and KGs, plus the last raw-text layer, which serves the purpose of backup knowledge in case the retrieval is not accomplished through the preliminary layers, formulate the precondition of the procedure of completion and condensation, thus can operate multi-layer query of PolyRAG framework.
III-B Knowledge Completion
As shown in the pyramid of Figure 1, the higher layers are more structured but not easy to define or extract, resulting in limited coverage. One common issue is the absence of important classes or attributes from the expert-defined Ontology schema, which has historically hindered the ease of defining Ontologies. By utilizing the pyramid framework, we can address this issue in a data-driven manner. The idea is to identify noteworthy concepts and relations that exist in the lower layers but are absent from the higher layers. These identified elements are then incorporated into the higher layers to enhance knowledge completion. For instance, our statistics indicate that while past_affiliations frequently appear in the knowledge graph and raw texts, but are missing from the Ontology. This oversight stems from the assumption in the initial Ontology that all facility members are affiliated with the same university, differing only by department, leading to the use of current_department. However, past_affiliations are valuable for user queries about alumni or consultants from other universities. Thus, it would be more effective to replace current_department with separate past_affiliations and current_affiliations to enhance the Ontology’s completeness and usability. These missing noteworthy concepts are indeed indicators for positions where the Ontology and knowledge graph differ significantly in semantic coverage. Therefore, if we can map the Ontology and knowledge graph into a same semantic space and those concepts can be identified by measuring the divergence of their coverage.
To achieve this, we begin by transforming class-attribute pairs in the Ontology layer $\mathcal{O}$ into subject+relation phrases that align with the format in the knowledge graph layer $\mathcal{K}$. Instructor embedding [55] of phrases at both layers are then encoded into a common semantic space. We then learn for each layer a multivariate Gaussian using